OpenStudy (anonymous):

Solve (z + 6)2 = 5 A. {-6+√5,6-√5} B. {-6±√5} C. {√5±-6} D. {6±√5}

OpenStudy (anonymous):

is the 2 supposed to be a squared?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay. Thanks

OpenStudy (australopithecus):

Rule you can multiple exponents on both sides of an equation so to solve for a: a^2 = b \[a^{2*(1/2)} = b^{1*(1/2)}\] \[a^{1} = b^{1/2}\] \[a = b^{1/2}\]

OpenStudy (australopithecus):

It might also help to remember BEDMAS order of operations.

OpenStudy (anonymous):

\[\sqrt{(Z+6)^{2}}=\sqrt{5}\] Then solve for Z.

OpenStudy (anonymous):

Would it be D?

OpenStudy (australopithecus):

I think you made a mistake in your algebra

OpenStudy (anonymous):

No, because if you subtract 6 from both sides the square root of 5 is not what is subtracted you subtract the 6.